OpenStudy (anonymous):

Help please :c If you have ever swum in a pool and your eyes began to sting and turn red, you are aware of the effects on an incorrect pH level. The pH level measures the concentration of hydronium ions and can be modeled by the function p(t) = –log10 t. The variable t represents the amount of hydronium ions, and p(t) gives the resulting pH level. Water at 25 degrees Celsius has a pH of 7. Anything that has a pH lower than 7 is called acidic, while pH's above 7 are basic or alkaline. Seawater has a pH just over 8, while lemonade has a pH of approximately 3.

OpenStudy (anonymous):

Create a graph of the pH function. Locate on your graph where the pH value is 0 and where it is 1. You may need to zoom in on your graph. I have the graph made I'm just not sure what the y values are.

OpenStudy (anonymous):

@phi @Hero

OpenStudy (anonymous):

@mathslover

hero (hero):

How did you make the graph without knowing what the y values are?

OpenStudy (anonymous):

I typed in the function they had being p(t)=-log10(t)

hero (hero):

If you typed y = -log(x), you'd get the same graph. Maybe that will help you.

OpenStudy (anonymous):

Okay, would I then do 0=-log(x)?

OpenStudy (anonymous):

My graph is different when I typed in what you told me to type. Maybe I just graphed wrong.

hero (hero):

When graphing, you usually begin with x and then figure out the corresponding y values.

hero (hero):

You probably graphed wrong.

hero (hero):

\(\log_{10}(x)\) is the same as \(\log(x)\)

OpenStudy (anonymous):

Okay. Anyways how would I begin to find the y values though? I have to find where the PH value is 0 and 1.

hero (hero):

If you have it graphed, you would see that -log(1) = 0 So (1,0) is definitely a point on the graph of y = -log(x) And I apologize, I meant for you to graph y = -log(x) (with the negative in front of the log).

OpenStudy (anonymous):

I'm sorry I'm terrible at math, and yeah I can see that (1,0) is a point

hero (hero):

And what does log(0) = ?

OpenStudy (anonymous):

Am I suppose to graph it or solve it on my calculator?

OpenStudy (anonymous):

I typed it in on Mathway and it says undefined

hero (hero):

Yeah but try -log(0.1)

OpenStudy (anonymous):

I got 1

hero (hero):

So that should help you out a bit. You will only be interested in values 0 < x < 1

OpenStudy (anonymous):

Okay.. Like I said I'm sorry and terrible with math.. I think I'm just going to call my teacher and see if he can help. Thanks a lot for everything and putting up with my idiocy! c:

hero (hero):

Have you tried log(0.000001) yet?

hero (hero):

That might help you with other PH Values

OpenStudy (anonymous):

No I haven't

OpenStudy (anonymous):

Okie dokie!

hero (hero):

-log(0.1) = 1 -log(0.01) = 2 -log(0.001) = 3 and so on..

hero (hero):

Anyway, good luck.